### v1

# Gluing diffeomorphisms, bi-Lipschitz mappings and homeomorphisms

Abstract.
In the paper, we address the following problem, often encountered in geometric constructions: having defined an orientation preserving diffeomorphism in a number of ‘patches’ on a manifold $\mathcal{M}$, can we extend it to a diffeomorphism of the whole manifold $\mathcal{M}$? We give a positive answer, provided the ‘patches’ are sufficiently regular. Then, we extend the same result to the case of bi-Lipschitz homeomorphisms and homeomorphisms.
Key words and phrases:
diffeomorphisms, homeomorphisms, bi-Lipschitz mappings, stable homeomorphisms, annulus theorem
2020 Mathematics Subject Classification:
Primary: 54C20 Secondary: 57N55, 58C07, 58C25
P.G. was supported by NCN grant no 2019/35/B/ST1/02030
Z.G. was supported by NCN grant no 2022/45/N/ST1/02977
P.H. was supported by NSF grant DMS-2055171
1. Introduction
[…]
Let $\mathcal{M}^{n}$ and $\mathcal{N}^{n}$ be two $n$-dimensional manifolds and let $\Omega\subset\mathcal{M}^{n}$ be open. By a homeomorphism $f:\Omega\to\mathcal{N}^{n}$ we mean a homeomorphism onto the image, i. e., a homeomorphism between $\Omega$ and $f(\Omega)$. Similar convention applies to diffeomorphisms.
If $A\subset\mathcal{M}^{n}$ is closed, then by a $C^{k}$-diffeomorphism $F:A\to\mathcal{N}^{n}$ we mean a mapping that extends to a $C^{k}$-diffeomorphism on an open neighborhood of $A$. Finally, a closed set $D\subset\mathcal{M}^{n}$ is a $C^{k}$-closed ball if there is a $C^{k}$-diffeomorphism $F:\mkern 1.7mu\overline{\mkern-1.7mu\mathbb{B}\mkern-1.5mu}\mkern 1.5mu^{n}\twoheadrightarrow D\subset\mathcal{M}^{n}$. In other words, if there is a $C^{k}$-diffeomorphism $F:B^{n}(0,1+\varepsilon)\to\mathcal{M}^{n}$ such that $F(\mkern 1.7mu\overline{\mkern-1.7muB\mkern-1.5mu}\mkern 1.5mu^{n}(0,1))=D$.
The aim of the paper is to prove Theorems 1.1, 1.6, 1.8 below.
Theorem 1.1.
Let $\mathcal{M}^{n}$ be an $n$-dimensional connected and oriented manifold of class $C^{k}$, $k\in\mathbb{N}\cup\{\infty\}$. Suppose that $\{{D}_{i}\}_{i=1}^{\ell}$ and $\{{D}^{\prime}_{i}\}_{i=1}^{\ell}$, ${D}_{i},{D}^{\prime}_{i}\subset\mathcal{M}^{n}$, are two families of pairwise disjoint $C^{k}$-closed balls. If $F_{i}:{D}_{i}\twoheadrightarrow{D}^{\prime}_{i}$, $i=1,2,\ldots,\ell$, are orientation preserving $C^{k}$-diffeomorphisms, then there is a $C^{k}$-diffeomorphism $F:\mathcal{M}^{n}\twoheadrightarrow\mathcal{M}^{n}$ such that $F|_{D_{i}}=F_{i}$. Moreover, if $D_{i}$ and $D^{\prime}_{i}$ for all $i=1,\ldots,\ell$ are contained in a domain $U\subset\mathcal{M}^{n}$, $F$ can be chosen to equal identity outside $U$.
Remark 1.2.
We do not assume that the balls in the family $\{{D}_{i}\}_{i=1}^{\ell}$ are disjoint from the balls in the family $\{{D}^{\prime}_{i}\}_{i=1}^{\ell}$.
Remark 1.3.
The diffeomorphisms $F_{i}$ can be very complicated and there is no apparent reason why an extension $F$ should exist: it follows from the celebrated theorem of Milnor [18, Theorem 5] that if $A\subset\mathbb{S}^{7}$ is a standard closed annulus, then there is an orientation preserving diffeomorphism $F:A\twoheadrightarrow A\subset\mathbb{S}^{7}$ that cannot be extended to a diffeomorphism of $\mathbb{S}^{7}$ (see Example 3.4).
Among the many possible corollaries based on Theorem 1.1, below we state one in the Euclidean setting, which is useful due to its elementary statement.
Corollary 1.4.
Let $\Omega\subset\mathbb{R}^{n}$ be open and let $D_{1}$ and $D_{2}$ be $C^{k}$-closed balls, $k\in\mathbb{N}\cup\{\infty\}$, such that $D_{2}\subset\mathring{D}_{1}\subset D_{1}\subset\Omega$. If $F:\Omega\to\mathbb{R}^{n}$ and $G:D_{2}\to\mathbb{R}^{n}$ are orientation preserving diffeomorphisms satisfying $G(D_{2})\subset F(\mathring{D}_{1})$, then there is a $C^{k}$-diffeomorphism $H:\Omega\twoheadrightarrow F(\Omega)$ that agrees with $F$ on $\Omega\setminus\mathring{D}_{1}$ and with $G$ on $D_{2}$.
Remark 1.5.
Let $D_{1}=\mkern 1.7mu\overline{\mkern-1.7muB\mkern-1.5mu}\mkern 1.5mu^{n}(0,2)$ and $D_{2}=\mkern 1.7mu\overline{\mkern-1.7muB\mkern-1.5mu}\mkern 1.5mu^{n}(0,1)$. If $F:\mathbb{R}^{n}\setminus\mathring{D}_{1}\twoheadrightarrow\mathbb{R}^{n}\setminus\mathring{D}_{1}$ is an orientation preserving diffeomorphism and $G=\operatorname{id}$ on $D_{2}$, one could expect a similar conclusion as in Corollary 1.4: existence of a diffeomorphism $H:\mathbb{R}^{n}\twoheadrightarrow\mathbb{R}^{n}$ that agrees with $F$ on $\mathbb{R}^{n}\setminus\mathring{D}_{1}$ and with $G$ on $D_{2}$. However, Milnor’s theorem, [18, Theorem 5], again yields a counterexample when $n=7$ (see Example 3.4).
We say that a closed set $D\subset\mathcal{M}^{n}$ is a bi-Lipschitz closed ball, if there is a bi-Lipschitz homeomorphism $F:\mkern 1.7mu\overline{\mkern-1.7mu\mathbb{B}\mkern-1.5mu}\mkern 1.5mu^{n}\twoheadrightarrow{D}$. We say that $D$ is a flat bi-Lipschitz closed ball if there is a bi-Lipschitz homeomorphism $F:\mkern 1.7mu\overline{\mkern-1.7mu\mathbb{B}\mkern-1.5mu}\mkern 1.5mu^{n}\twoheadrightarrow{D}$ that can be extended to a bi-Lipschitz map on a neighborhood of $\mkern 1.7mu\overline{\mkern-1.7mu\mathbb{B}\mkern-1.5mu}\mkern 1.5mu^{n}$. Not every bi-Lipschitz ball is flat, as the example of the bi-Lipschitz Fox-Artin ball shows [7], [8, Theorem 7.2.1], [17, Theorem 3.7]. For that reason in the next result we need to assume flatness of the balls.
Theorem 1.6.
Let $\mathcal{M}^{n}$ be an $n$-dimensional connected and oriented Lipschitz manifold. Suppose that $\{{D}_{i}\}_{i=1}^{\ell}$ and $\{{D}^{\prime}_{i}\}_{i=1}^{\ell}$, ${D}_{i},{D}^{\prime}_{i}\subset\mathcal{M}^{n}$, are two families of pairwise disjoint flat bi-Lipschitz closed balls. If $F_{i}:{D}_{i}\twoheadrightarrow{D}^{\prime}_{i}$, $i=1,2,\ldots,\ell$ are orientation preserving bi-Lipschitz homeomorphisms, then there is a bi-Lipschitz homeomorphism $F:\mathcal{M}^{n}\twoheadrightarrow\mathcal{M}^{n}$ such that $F|_{D_{i}}=F_{i}$. Moreover, if $D_{i}$ and $D^{\prime}_{i}$ for all $i=1,\ldots,\ell$ are contained in a domain $U\subset\mathcal{M}^{n}$, $F$ can be chosen to equal identity outside $U$.
Remark 1.7.
Let us note that it follows from the Schönflies theorem for Lipschitz maps [26, Theorem A] that on the plane every bi-Lipschitz closed ball is flat, and hence the assumption on bi-Lipschitz flatness is superfluous for $n=2$, see also [6, 15] for recent related results. For a geometric characterization of flat bi-Lipschitz balls in $\mathbb{R}^{n}$, $n\geq 2$, see Remark 4.6.
While the overall strategy for Theorem 1.6 is similar to that for Theorem 1.1, the problem is much deeper, because it requires the bi-Lipschitz stable homeomorphism theorem proved by Sullivan [25] (see Lemma 4.3 below). This is a very difficult result, but in dimensions $n\leq 3$ there is a simpler proof [26], [28, Theorem 3.2]. Unfortunately, the proof in the case $n\geq 4$ lacks details leading to some doubts, see [27, Introduction] and [11].
We say that a closed set $D\subset\mathcal{M}^{n}$ is a topological closed ball, if there is a homeomorphism $F:\mkern 1.7mu\overline{\mkern-1.7mu\mathbb{B}\mkern-1.5mu}\mkern 1.5mu^{n}\twoheadrightarrow{D}$. We say that $D$ is a flat topological closed ball if there is a homeomorphism $F:\mkern 1.7mu\overline{\mkern-1.7mu\mathbb{B}\mkern-1.5mu}\mkern 1.5mu^{n}\twoheadrightarrow{D}$ that can be extended to a homeomorphism on a neighborhood of $\mkern 1.7mu\overline{\mkern-1.7mu\mathbb{B}\mkern-1.5mu}\mkern 1.5mu^{n}$ (see also Remark 4.6). Examples of wild spheres (the Alexander horned sphere) show that not every closed topological ball is flat, so, as in the bi-Lipschitz case, we need to assume flatness of balls.
Theorem 1.8.
Let $\mathcal{M}^{n}$ be an $n$-dimensional connected and oriented topological manifold. Suppose that $\{{D}_{i}\}_{i=1}^{\ell}$ and $\{{D}^{\prime}_{i}\}_{i=1}^{\ell}$, ${D}_{i},{D}^{\prime}_{i}\subset\mathcal{M}^{n}$, are two families of pairwise disjoint flat topological closed balls. If $F_{i}:{D}_{i}\twoheadrightarrow{D}^{\prime}_{i}$, $i=1,2,\ldots,\ell$ are orientation preserving homeomorphisms, then there is a homeomorphism $F:\mathcal{M}^{n}\twoheadrightarrow\mathcal{M}^{n}$ such that $F|_{D_{i}}=F_{i}$. Moreover, if $D_{i}$ and $D^{\prime}_{i}$ for all $i=1,\ldots,\ell$ are contained in a domain $U\subset\mathcal{M}^{n}$, $F$ can be chosen to equal identity outside $U$.
The proof of this result requires the stable homeomorphism theorem, see Lemma 4.3.
In such topological problems as Theorems 1.6 and 1.8, the dimension $n=4$ is often doubtful, but the results are also true for $n=4$.
Below we state an analogue of Corollary 1.4 in the bi-Lipschitz and in the homeomorphic case.
Corollary 1.9.
Let $\Omega\subset\mathbb{R}^{n}$ be open and let $D_{1}$ and $D_{2}$ be flat topological (or flat bi-Lipschitz) closed balls such that $D_{2}\subset\mathring{D}_{1}\subset D_{1}\subset\Omega$. If $F:\Omega\to\mathbb{R}^{n}$ and $G:D_{2}\to\mathbb{R}^{n}$ are orientation preserving (bi-Lipschitz) homeomorphisms satisfying $G(D_{2})\subset F(\mathring{D}_{1})$, then there is a (bi-Lipschitz) homeomorphism $H:\Omega\twoheadrightarrow F(\Omega)$ that agrees with $F$ on $\Omega\setminus\mathring{D}_{1}$ and with $G$ on $D_{2}$.
Theorem 1.1 for $\ell=1$ was proved by Palais [21]. In that case the result is about extending diffeomorphisms, while the case $\ell\geq 2$ is about gluing them. Our proof is based on Palais’ argument (see Lemma 2.2). More importantly, we show that the same method can be extended to the case of homeomorphisms and bi-Lipschitz homeomorphisms. While some readers may view the above results, especially Theorem 1.1, as folklore, surprisingly, we could not locate such results in the literature. The results appear natural and useful, and we believe they warrant detailed and easy-to-follow proofs. The results might be useful to mathematicians who are not familiar with advanced geometric topology and that was one of our motivation in writing the paper.
The paper is structured as follows. In Section 2 we explain a beautiful argument of Palais [21]. In Section 3 we prove Theorem 1.1 and Corollary 1.4. The remaining sections are devoted to the proofs of Theorems 1.6 and 1.8. While the overall strategy is similar to that of the proof of Theorem 1.1, the main difficulty is in the generalization of Lemma 2.1. This lemma is easy in the smooth case, but very difficult in the case of homeomorphisms and bi-Lipschitz homeomorphisms. In Section 4 we collect known facts about the stable homeomorphism theorem and the annulus theorem in both cases, of homeomorphisms and bi-Lipschitz homeomorphisms. At the end of Section 4 we prove Lemma 4.9, which is a counterpart of Lemma 2.1. After that, the proofs of Theorems 1.6 and 1.8 are almost the same as the proof of Theorem 1.1 and we show these proofs in Section 5.
Acknowledgments
We would like to thank Igor Belegradek for directing us to the paper [21] of Palais.
2. Palais
A well known result [20, Lemma 8.1], [22, Theorem 5.5], says that any orientation preserving $C^{k}$-diffeomorphism that fixes the origin, can be linearized to identity on a sufficiently small ball.
Lemma 2.1.
Suppose that $H:\mkern 1.7mu\overline{\mkern-1.7muB\mkern-1.5mu}\mkern 1.5mu(0,\varrho)\to\mathbb{R}^{n}$ is an orientation preserving $C^{k}$-diffeomorphism, $k\in\mathbb{N}\cup\{\infty\}$, with $H(0)=0$. Then there is a $\delta\in(0,\varrho/2)$ and $H_{1}:\mkern 1.7mu\overline{\mkern-1.7muB\mkern-1.5mu}\mkern 1.5mu(0,\varrho)\to\mathbb{R}^{n}$, $C^{k}$-diffeomorphism, such that $H_{1}=H$ on $\mkern 1.7mu\overline{\mkern-1.7muB\mkern-1.5mu}\mkern 1.5mu(0,\varrho)\setminus B(0,\varrho/2)$ and $H_{1}=\operatorname{id}$ on $B(0,\delta)$.
The idea is to glue $H$ with the linear mapping $DH(0)$ defined in a small neighborhood of 00 using a suitable partition of unity and then to deform the linear map to the identity on a smaller neighborhood of 00 using the fact that $DH(0)$ and the identity belong to the same connected component of $GL(n)$. One also needs some topology to ensure that the resulting mapping is a diffeomorphism and not only a local diffeomorphism. For a detailed proof, we refer to Lemma 3.8 and Lemma 3.10 from [10].
The next surprising result was proved by Palais [21]. The proof is short, elementary, beautiful, and unexpected. It is the core of the proof of Theorem 1.1.
Lemma 2.2.
Suppose that $H:\mkern 1.7mu\overline{\mkern-1.7muB\mkern-1.5mu}\mkern 1.5mu(0,\varrho)\to\mathbb{R}^{n}$ is an orientation preserving $C^{k}$-diffeomorphism, $k\in\mathbb{N}\cup\{\infty\}$, with $H(0)=0$. Then for any $\varepsilon>0$, there is a $C^{k}$-diffeomorphism $\widetilde{H}:\mathbb{R}^{n}\twoheadrightarrow\mathbb{R}^{n}$ such that
(2.1) $\widetilde{H}(x)=\begin{cases}H(x)&\text{ if }x\in\mkern 1.7mu\overline{\mkern-1.7muB\mkern-1.5mu}\mkern 1.5mu(0,\varrho),\\ x&\text{ if }\operatorname{dist}(x,A)\geq\varepsilon,\end{cases}$
where $A=\mkern 1.7mu\overline{\mkern-1.7muB\mkern-1.5mu}\mkern 1.5mu(0,\varrho)\cup H(\mkern 1.7mu\overline{\mkern-1.7muB\mkern-1.5mu}\mkern 1.5mu(0,\varrho))$.
Remark 2.3.
If $\widetilde{H}:\mathbb{R}^{n}\twoheadrightarrow\mathbb{R}^{n}$ is a diffeomorphism such that $\widetilde{H}(0)=0$, but $\widetilde{H}(x)\neq x$ for $0\neq x\in\mkern 1.7mu\overline{\mkern-1.7muB\mkern-1.5mu}\mkern 1.5mu(0,\varrho)$, then clearly $\widetilde{H}(x)\neq x$ for $0\neq x\in A$ and hence $\widetilde{H}(x)\neq x$ in a neighborhood of $\partial A$. Therefore, the condition $\widetilde{H}(x)=x$ if $\operatorname{dist}(x,A)\geq\varepsilon$ in (2.1) is sharp.
Proof.
Let $\varepsilon>0$ be given. By assumption, $H$ extends to a $C^{k}$-diffeomorphism on $B(0,\varrho+3\tau)$ for some $\tau>0$; we denote this extension by $H$ as well. By decreasing $\tau>0$ if necessary, we may assume that
$B(0,\varrho+3\tau)\cup H(B(0,\varrho+3\tau))\subset\{x:\,\operatorname{dist}(x,A)<\varepsilon\}.$
By Lemma 2.1, there is a $\delta\in(0,\varrho/2)$ and a $C^{k}$-diffeomorphism $H_{1}:B(0,\varrho+3\tau)\to\mathbb{R}^{n}$ such that
$H_{1}(x)=\begin{cases}H(x)&\text{ if }x\in B(0,\varrho+3\tau)\setminus B(0,\varrho/2),\\ x&\text{ if }x\in B(0,\delta).\end{cases}$
Let $\phi:\mathbb{R}\to\mathbb{R}$ be a non-decreasing, smooth function satisfying
$\phi(t)=\begin{cases}1&\text{ if }t>\varrho+2\tau,\\ \delta(\varrho+\tau)^{-1}&\text{ if }t<\varrho+\tau,\end{cases}$
and define $\Phi:\mathbb{R}^{n}\twoheadrightarrow\mathbb{R}^{n}$ by $\Phi(x):=\phi(|x|)x$. The mapping $\Phi$ is obviously a smooth diffeomorphism of $\mathbb{R}^{n}$. Note also that $\Phi(x)$ on $B(0,\varrho+\tau)$ acts as scaling by a factor of $\delta(\varrho+\tau)^{-1}$, $\Phi(B(0,\varrho+\tau))=B(0,\delta)$, while $\Phi(x)=x$ when $|x|>\varrho+2\tau$. Since $B(0,\varrho+2\tau)\subset\{\operatorname{dist}(x,A)<\varepsilon\}$, $\Phi=\operatorname{id}$ in $\{\operatorname{dist}(x,A)\geq\varepsilon\}$.
Consider the $C^{k}$-diffeomorphism
(2.2) $H_{1}\circ\Phi^{-1}\circ H_{1}^{-1}:H_{1}(B(0,\varrho+3\tau))\to\mathbb{R}^{n}.$
It is a well defined diffeomorphism, because $\Phi$ maps the ball $B(0,\varrho+3\tau)$ onto itself. The diffeomorphism defined in (2.2) is identity near the boundary of $H_{1}(B(0,\varrho+3\tau))$, because
(2.3) $H_{1}\circ\Phi^{-1}\circ H_{1}^{-1}(x)=x\quad\text{for}\quad x\in H_{1}\big{(}B(0,\varrho+3\tau)\setminus B(0,\varrho+2\tau)\big{)}.$
Indeed, $\Phi^{-1}$ is the identity on $B(0,\varrho+3\tau)\setminus B(0,\varrho+2\tau)$, so for $x$ as in (2.3), we have $\Phi^{-1}(H_{1}^{-1}(x))=H_{1}^{-1}(x)$. Therefore, the $C^{k}$-diffeomorphism (2.2) has the extension to a $C^{k}$-diffeomorphism of $\mathbb{R}^{n}$ by identity:
[…]
Moreover, the mapping defined in (2.4) is identity in $\{x:\,\operatorname{dist}(x,A)\geq\varepsilon\}$, because $H_{1}(B(0,\varrho+3\tau))=H(B(0,\varrho+3\tau))\subset\{\operatorname{dist}(x,A)<\varepsilon\}$.
Now, we define the $C^{k}$-diffeomorphism
[…]
Note that since $\Phi$ and the diffeomorphism defined in (2.4) are both equal to identity in the set $\{\operatorname{dist}(x,A)\geq\varepsilon\}$, so is $H_{2}$.
If $x\in B(0,\varrho+\tau)$, then
$\Phi(x)=\delta(\varrho+\tau)^{-1}x\in B(0,\delta)\subset H_{1}(B(0,\varrho+3\tau)).$
Thus by (2.4), and taking into account that $H_{1}^{-1}=\operatorname{id}$ on $B(0,\delta)$, for $x\in B(0,\varrho+\tau)$ we have
$H_{2}(x)=H_{1}\circ\Phi^{-1}\circ H_{1}^{-1}(\Phi(x))=H_{1}\circ\Phi^{-1}\circ\Phi(x)=H_{1}(x).$
Finally, since $H_{2}=H_{1}=H$ near $\partial B(0,\varrho)$, we can set
[…]
and, clearly, $\widetilde{H}$ is a diffeomorphism that satisfies (2.1). ∎
Corollary 2.4.
Let $G:\mkern 1.7mu\overline{\mkern-1.7muB\mkern-1.5mu}\mkern 1.5mu(a,r)\twoheadrightarrow\mkern 1.7mu\overline{\mkern-1.7muB\mkern-1.5mu}\mkern 1.5mu(a,r)$ be an orientation preserving $C^{k}$-diffeomorphism, $k\in\mathbb{N}\cup\{\infty\}$. Then for any $\varepsilon>0$, there is a $C^{k}$-diffeomorphism $\widetilde{G}:\mathbb{R}^{n}\twoheadrightarrow\mathbb{R}^{n}$ such that $\widetilde{G}=G$ on $\mkern 1.7mu\overline{\mkern-1.7muB\mkern-1.5mu}\mkern 1.5mu(a,r)$ and $\widetilde{G}=\operatorname{id}$ outside $B(a,r+\varepsilon)$.
Proof.
If $G(a)=a$, it is a straightforward consequence of Lemma 2.2. If not, we construct a $C^{\infty}$-diffeomorphism $F:\mathbb{R}^{n}\twoheadrightarrow\mathbb{R}^{n}$ such that $F(G(a))=a$ and $F=\operatorname{id}$ in $\mathbb{R}^{n}\setminus B(a,r)$. One can construct $F$ through a $1$-parameter group of diffeomorphisms generated by a compactly supported vector field.
Let $T(x)=x-a$ and $G_{1}=T\circ F\circ G\circ T^{-1}$. Then $G_{1}$ is an orientation preserving $C^{k}$-diffeomorphism of $\mkern 1.7mu\overline{\mkern-1.7muB\mkern-1.5mu}\mkern 1.5mu(0,r)$ onto itself, $G_{1}(0)=0$. For any $\varepsilon>0$, we may thus apply Lemma 2.2 to $G_{1}$, obtaining a diffeomorphism $\widetilde{G}_{1}:\mathbb{R}^{n}\twoheadrightarrow\mathbb{R}^{n}$ such that
$\widetilde{G}_{1}(x)=\begin{cases}G_{1}(x)&\text{ for }x\in\mkern 1.7mu\overline{\mkern-1.7muB\mkern-1.5mu}\mkern 1.5mu(0,r),\\ x&\text{ for }x\not\in B(0,r+\varepsilon).\end{cases}$
Now, $\widetilde{G}=F^{-1}\circ T^{-1}\circ\widetilde{G}_{1}\circ T$ satisfies the claim of the lemma. ∎
3. Proof of Theorem 1.1
Lemma 3.1.
Let $\{D_{i}\}_{i=1}^{\ell}$, $D_{i}\subset\mathbb{R}^{n}$, be a family of pairwise disjoint $C^{k}$-closed balls and let $p_{i}\in\mathring{D}_{i}$, $i=1,\ldots,\ell$, be given. Then there is $\varepsilon_{o}>0$ such that for any $\varepsilon\in(0,\varepsilon_{o})$ there is a $C^{k}$-diffeomorphism $F_{\varepsilon}:\mathbb{R}^{n}\twoheadrightarrow\mathbb{R}^{n}$ satisfying $F_{\varepsilon}(\mkern 1.7mu\overline{\mkern-1.7muB\mkern-1.5mu}\mkern 1.5mu(p_{i},\varepsilon))=D_{i}$ for $i=1,\ldots,\ell$ and
$F_{\varepsilon}(x)=x\text{ if }\operatorname{dist}\Big{(}x,\bigcup\nolimits_{i=1}^{\ell}D_{i}\Big{)}\geq\varepsilon.$
Proof.
Let $0<\varepsilon_{o}\leq\frac{1}{4}\min_{i\neq j}\operatorname{dist}(D_{i},D_{j})$ be such that $\mkern 1.7mu\overline{\mkern-1.7muB\mkern-1.5mu}\mkern 1.5mu(p_{i},\varepsilon_{o})\subset\mathring{D}_{i}$ for all $i=1,2,\ldots,\ell$ and let $\varepsilon\in(0,\varepsilon_{o})$. By assumption, there are $C^{k}$-diffeomorphisms $H_{i}:\mkern 1.7mu\overline{\mkern-1.7muB\mkern-1.5mu}\mkern 1.5mu(p_{i},\varepsilon)\twoheadrightarrow D_{i}$. Additionally, we can assume that $H_{i}(p_{i})=p_{i}$. Indeed, if $H_{i}(p_{i})\neq p_{i}$, we compose $H_{i}$ with a diffeomorphism of $D_{i}$ which is identity near $\partial D_{i}$ and maps $H_{i}(p_{i})$ to $p_{i}$. Such a diffeomorphism can be constructed through a $1$-parameter group of diffeomorphisms generated by a compactly supported vector field. Note that
$\mkern 1.7mu\overline{\mkern-1.7muB\mkern-1.5mu}\mkern 1.5mu(p_{i},\varepsilon)\cup H_{i}(\mkern 1.7mu\overline{\mkern-1.7muB\mkern-1.5mu}\mkern 1.5mu(p_{i},\varepsilon))=D_{i}\quad\text{for }i=1,2,\ldots,\ell.$
According to Lemma 2.2, $H_{i}$ can be extended to a diffeomorphism $\widetilde{H}_{i}:\mathbb{R}^{n}\twoheadrightarrow\mathbb{R}^{n}$ that is identity outside the $\varepsilon$-neighborhood of $D_{i}$. Since the sets $\{\operatorname{dist}(x,D_{i})\leq\varepsilon\}$ are pairwise disjoint, we can glue the diffeomorphisms $\widetilde{H}_{i}$ to a diffeomorphism $F_{\varepsilon}$ satisfying the claim of the lemma. ∎
The next result is well known. The diffeomorphism $H$ can be constructed using $1$-parameter groups of diffeomorphisms with compact support in $U$, see e.g. [10, Lemma 3.13].
Lemma 3.2.
Let $U\subset\mathbb{R}^{n}$ be a domain and let $\{p_{i}\}_{i=1}^{\ell}$ and $\{q_{i}\}_{i=1}^{\ell}$ be given points in $U$such that $p_{i}\neq p_{j}$ and $q_{i}\neq q_{j}$ for $i\neq j$. Then, there exists an $\varepsilon_{o}>0$ such that for any $\varepsilon\in(0,\varepsilon_{o})$ there is a $C^{\infty}$-diffeomorphism $H:\mathbb{R}^{n}\twoheadrightarrow\mathbb{R}^{n}$, identity on $\mathbb{R}^{n}\setminus U$, such that
(3.1) $H(\mkern 1.7mu\overline{\mkern-1.7muB\mkern-1.5mu}\mkern 1.5mu(q_{i},\varepsilon))=\mkern 1.7mu\overline{\mkern-1.7muB\mkern-1.5mu}\mkern 1.5mu(p_{i},\varepsilon).$
Combining Lemmata 3.1 and 3.2 we obtain the following result, which is a Euclidean variant of Theorem 1.1.
Corollary 3.3.
Let $U\subset\mathbb{R}^{n}$ be a domain. Suppose that $\{D_{i}\}_{i=1}^{\ell}$ and $\{D^{\prime}_{i}\}_{i=1}^{\ell}$, $D_{i},D^{\prime}_{i}\subset U$, are two families of pairwise disjoint $C^{k}$-closed balls. If $F_{i}:D_{i}\twoheadrightarrow D^{\prime}_{i}$, $i=1,2,\ldots,\ell$, are orientation preserving $C^{k}$-diffeomorphisms, then there is a $C^{k}$-diffeomorphism ${F:\mathbb{R}^{n}\twoheadrightarrow\mathbb{R}^{n}}$ such that $F|_{D_{i}}=F_{i}$ and $F=\operatorname{id}$ in $\mathbb{R}^{n}\setminus U$.
Proof.
Let $p_{i}\in\mathring{D}_{i}$ and $q_{i}\in\mathring{D}_{i}^{\prime}$. By Lemma 3.1, there are diffeomorphisms ${A_{1},A_{2}:\mathbb{R}^{n}\twoheadrightarrow\mathbb{R}^{n}}$ such that
$A_{1}(\mkern 1.7mu\overline{\mkern-1.7muB\mkern-1.5mu}\mkern 1.5mu(p_{i},\varepsilon))=D_{i},\quad A_{2}(\mkern 1.7mu\overline{\mkern-1.7muB\mkern-1.5mu}\mkern 1.5mu(q_{i},\varepsilon))=D_{i}^{\prime},\quad\text{and}\quad A_{1,2}=\operatorname{id}\text{ in }\mathbb{R}^{n}\setminus U.$
This is true provided $\varepsilon>0$ is sufficiently small. In what follows we will not clarify how small $\varepsilon$ should be, but it will always be clear from the context.
Let $H:\mathbb{R}^{n}\twoheadrightarrow\mathbb{R}^{n}$ be a diffeomorphism as in Lemma 3.2. Then,
$G_{i}:=H\circ A_{2}^{-1}\circ F_{i}\circ A_{1}:\mkern 1.7mu\overline{\mkern-1.7muB\mkern-1.5mu}\mkern 1.5mu(p_{i},\varepsilon)\twoheadrightarrow\mkern 1.7mu\overline{\mkern-1.7muB\mkern-1.5mu}\mkern 1.5mu(p_{i},\varepsilon),\quad i=1,2,\ldots,\ell$
are orientation preserving diffeomorphisms. According to Corollary 2.4 each of them can be extended to a diffeomorphism $\widetilde{G}_{i}:\mathbb{R}^{n}\twoheadrightarrow\mathbb{R}^{n}$ that is identity outside an arbitrarily small neighborhood of $\mkern 1.7mu\overline{\mkern-1.7muB\mkern-1.5mu}\mkern 1.5mu(p_{i},\varepsilon)$ so they can be glued to a diffeomorphism $G:\mathbb{R}^{n}\twoheadrightarrow\mathbb{R}^{n}$ satisfying $G|_{\mkern 1.7mu\overline{\mkern-1.7muB\mkern-1.5mu}\mkern 1.5mu(p_{i},\varepsilon)}=G_{i}$. We can clearly guarantee that $G=\operatorname{id}$ in $\mathbb{R}^{n}\setminus U$. Then, it easily follows that $F:=A_{2}\circ H^{-1}\circ G\circ A_{1}^{-1}$ satisfies the claim of the corollary. ∎
Proof of Corollary 1.4.
Since $F^{-1}\circ G:D_{2}\to\mathring{D}_{1}$ is an orientation preserving diffeomorphism between $C^{k}$-closed balls $D_{2},(F^{-1}\circ G)(D_{2})\subset\mathring{D}_{1}$, it follows from Corollary 3.3 that there is a diffeomorphism $\Phi:\mathbb{R}^{n}\twoheadrightarrow\mathbb{R}^{n}$ such that $\Phi=F^{-1}\circ G$ in $D_{2}$ and $\Phi=\operatorname{id}$ in $\mathbb{R}^{n}\setminus\mathring{D}_{1}$. Then the diffeomorphism $H:=F\circ\Phi:\Omega\to\mathbb{R}^{n}$ is well defined and satisfies $H|_{D_{2}}=G$, $H|_{\Omega\setminus\mathring{D}_{1}}=F$. ∎
Proof of Theorem 1.1..
We prove the result by reducing it to the Euclidean setting, i.e., to the case treated in Corollary 3.3.
Let $U$ be a domain containing $D_{i},D^{\prime}_{i}$ for $i=1,\ldots,\ell$ (note that we include the case $U=\mathcal{M}^{n}$). Let $K\subset U$ be a $C^{k}$-closed ball. Then, there are $C^{k}$-diffeomorphisms $H,H^{\prime}:\mathcal{M}^{n}\twoheadrightarrow\mathcal{M}^{n}$ such that $H(D_{i})\subset\mathring{K}$, $H^{\prime}(D_{i}^{\prime})\subset\mathring{K}$ for all $i=1,2,\ldots,\ell$, and $H$ and $H^{\prime}$ are identity outside $U$.
We will sketch the construction of $H$ leaving details to the reader. The construction of $H^{\prime}$ is analogous. Since $D_{i}$ is diffeomorphic to $\mkern 1.7mu\overline{\mkern-1.7mu\mathbb{B}\mkern-1.5mu}\mkern 1.5mu^{n}$, we can find a diffeomorphism $\Psi_{i}:\mathcal{M}^{n}\twoheadrightarrow\mathcal{M}^{n}$, identity outside a small neighborhood of $D_{i}$, that maps $D_{i}$ into a small neighborhood of $p_{i}\in\mathring{D}_{i}$ (cf. the construction of $\Phi$ in the proof of Lemma 2.2). Since the sets $D_{i}$ are pairwise disjoint, we can glue the diffeomorphisms $\Psi_{i}$ to $\Psi:\mathcal{M}^{n}\twoheadrightarrow\mathcal{M}$ that is identity outside a small neighborhood of $\bigcup_{i=1}^{\ell}D_{i}$, and hence identity outside $U$.
Next, we construct a vector field $X$ with compact support in $U$, whose $1$-parameter group of diffeomorphisms $\Phi_{t}$ satisfies $\Phi_{t_{o}}(\Psi(D_{i}))\subset\mathring{K}$ for all $i$, provided $t_{o}>0$ is sufficiently large. To construct such a vector field, we take smooth arcs $\gamma_{i}$ in $U$, connecting $p_{i}$ to some $p_{i}^{\prime}\in\mathring{K}$ and making sure that $\gamma_{i}\cap\gamma_{j}=\varnothing$ if $i\neq j$. Then, we take vector fields along $\gamma_{i}$ in the direction from $p_{i}$ to $p_{i}^{\prime}$, and extend them to compactly supported vector fields $X_{i}$ on $\mathcal{M}^{n}$ with the supports near $\gamma_{i}$. Clearly, we can guarantee that $\operatorname{supp}X_{i}\cap\operatorname{supp}X_{j}=\varnothing$ if $i\neq j$, and $\operatorname{supp}X_{i}\subset U$ for all $i$, so the vector fields glue to a compactly supported vector field $X$ on $\mathcal{M}^{n}$ with $\operatorname{supp}X\subset U$. Let $\Phi_{t}$ be the 1-parameter group of diffeomorphisms generated by $X$. If $\Psi(D_{i})$ lies in a sufficiently small neighborhood of $p_{i}$ (which we can guarantee), then $\Phi_{t_{o}}(\Psi(D_{i}))\subset\mathring{K}$ for all $i$ and sufficiently large $t_{o}>0$. Thus, we can take $H:=\Phi_{t_{o}}\circ\Psi$. Clearly, $H=\operatorname{id}$ in $\mathcal{M}^{n}\setminus U$.
Let $\widehat{F}_{i}:H(D_{i})\twoheadrightarrow H^{\prime}(D_{i}^{\prime})$, $\widehat{F}_{i}=H^{\prime}\circ F_{i}\circ H^{-1}$. If we represent $K$ in a coordinate system, Corollary 3.3 allows us to construct a diffeomorphism $\widehat{F}:K\twoheadrightarrow K$ such that
$\widehat{F}|_{H(D_{i})}=\widehat{F}_{i}\quad\text{and}\quad\widehat{F}=\operatorname{id}\text{ near }\partial K,$
so $\widehat{F}$ extends to a diffeomorphism $\widehat{F}:\mathcal{M}^{n}\twoheadrightarrow\mathcal{M}^{n}$ by identity outside $K$. Now, it easily follows that
$F:=(H^{\prime})^{-1}\circ\widehat{F}\circ H:\mathcal{M}^{n}\twoheadrightarrow\mathcal{M}^{n}$
satisfies $F|_{D_{i}}=F_{i}$ and $F=\operatorname{id}$ in $\mathcal{M}^{n}\setminus U$. ∎
Example 3.4.
In his celebrated article [18] introducing exotic $7$-spheres, Milnor proved the existence of an orientation preserving $C^{\infty}$-diffeomorphism $f:\mathbb{S}^{6}\twoheadrightarrow\mathbb{S}^{6}$ which is not isotopic through diffeomorphisms to identity ([18, Theorem 5]). This diffeomorphism can be extended radially as a $C^{\infty}$-diffeomorphism $\mathbb{R}^{7}\setminus\{0\}\to\mathbb{R}^{7}$. However, it cannot be extended to a $C^{\infty}$-diffeomorphism onto $\mathbb{R}^{7}\twoheadrightarrow\mathbb{R}^{7}$. If it were, it would be possible to find a $C^{\infty}$-diffeomorphism $G$ which coincides with $f$ on $\mathbb{S}^{6}$ and which is equal to identity on a small ball $B(0,r)$, like in Lemma 2.1. Since the annulus $\mkern 1.7mu\overline{\mkern-1.7mu\mathbb{B}\mkern-1.5mu}\mkern 1.5mu^{7}\setminus B(0,r)$ is diffeomorphic to $\mathbb{S}^{6}\times[0,1]$, we would have a diffeomorphism $F:\mathbb{S}^{6}\times[0,1]\twoheadrightarrow\mathbb{S}^{6}\times[0,1]$, $F(x,1)=f(x)$, $F(x,0)=x$. Such a diffeomorphism need not be an isotopy (the latter must preserve the sets $\mathbb{S}^{6}\times\{t\}$), but it is a pseudoisotopy between $f$ and the identity, and since we discuss diffeomorphisms of $\mathbb{S}^{6}$, by Cerf’s pseudoisotopy-to-isotopy theorem ([4], see also [13, Theorem 2.71]) this implies that $f$ is isotopic to the identity, which is a contradiction.
Also, if we consider $\mathbb{S}^{6}$ to be embedded into $\mathbb{S}^{7}$, we see that the same $f$ can be extended to a $C^{\infty}$-homeomorphism on an annulus $A\subset\mathbb{S}^{7}$, but cannot be extended to a $C^{\infty}$-diffeomorphism of $\mathbb{S}^{7}$.
4. Topological local linearization
Definition 4.1.
Let $\varrho>0$. We say that a homeomorphism $F:\mathbb{R}^{n}\twoheadrightarrow\mathbb{R}^{n}$ (or $F:B(0,\varrho)\twoheadrightarrow B(0,\varrho)$) is stable if
(4.1) $F=f_{1}\circ\ldots\circ f_{k}$
for some $k\in\mathbb{N}$ and homeomorphisms $f_{i}:\mathbb{R}^{n}\twoheadrightarrow\mathbb{R}^{n}$ ($f_{i}:B(0,\varrho)\twoheadrightarrow B(0,\varrho)$) such that $f_{i}|_{U_{i}}=\operatorname{id}$ for some nonempty open set $U_{i}\subset\mathbb{R}^{n}$ ($U_{i}\subset B(0,\varrho)$).
We say that a homeomorphism $F:\mathbb{R}^{n}\twoheadrightarrow\mathbb{R}^{n}$ (or $F:B(0,\varrho)\twoheadrightarrow B(0,\varrho)$) is locally bi-Lipschitz stable if $F$ can be written as in (4.1) for some $k\in\mathbb{N}$ and for locally bi-Lipschitz homeomorphisms $f_{i}:\mathbb{R}^{n}\twoheadrightarrow\mathbb{R}^{n}$ ($f_{i}:B(0,\varrho)\twoheadrightarrow B(0,\varrho)$) such that $f_{i}|_{U_{i}}=\operatorname{id}$ for some nonempty open set $U_{i}\subset\mathbb{R}^{n}$ ($U_{i}\subset B(0,\varrho)$).
Remark 4.2.
Without loss of generality, we may assume that the $U_{i}$ in the definition above are pairwise disjoint balls with the same radius (compactly contained in $B(0,\varrho)$).
Lemma 4.3 (Stable homeomorphism theorem).
Any orientation preserving homeomorphism $F:\mathbb{R}^{n}\twoheadrightarrow\mathbb{R}^{n}$ is stable. If $F$ is additionally locally bi-Lipschitz, then $F$ is locally bi-Lipschitz stable.
The stable homeomorphism theorem in the homeomorphic case in dimensions $n=2,3$ is a classical result due to Radó [24] for $n=2$ and Moise [19] for $n=3$. For $n>4$, it was proved by Kirby in [14] and for $n=4$ by Quinn in [23] (see page 1 and Theorem 2.2.2). For a short explanation of the intricacies of the proof in the topological case, we refer to [12]. In the bi-Lipschitz case, the stable homeomorphism theorem was derived in [27, Theorem 3.12] from the existence of groups known as Sullivan groups, whose existence was claimed by Sullivan in [25].
Since the open unit ball is homeomorphic to $\mathbb{R}^{n}$ and this homeomorphism can be chosen to be locally bi-Lipschitz, the following is easy to prove.
Corollary 4.4.
An orientation preserving homeomorphism $F:B(0,\varrho)\twoheadrightarrow B(0,\varrho)$ is stable. If, additionally, $F$ is locally bi-Lipschitz, then $F$ is also locally bi-Lipschitz stable.
The next result is a version of Lemma 2.1 in the case of homeomorphisms and bi-Lipschitz maps. However, it is not a good one, because the assumption that $F$ maps $B(0,\varrho)$ onto itself is very restrictive. We will remove this assumption later in Lemma 4.9 using the annulus theorem.
Lemma 4.5.
Let $F:B(0,\varrho)\twoheadrightarrow B(0,\varrho)$ be an orientation preserving homeomorphism. Then for any $\delta\in(0,\varrho)$, there is $\tau\in(\delta,\varrho)$ and a homeomorphism $\widehat{F}:B(0,\varrho)\twoheadrightarrow B(0,\varrho)$ such that
(4.2) $\widehat{F}(x)=\ \begin{cases}F(x)&\text{for }x\in B(0,\varrho)\setminus B(0,\tau),\\ x&\text{ for }x\in B(0,\delta).\end{cases}$
If $F$ is locally bi-Lipschitz, then $\widehat{F}$ is locally bi-Lipschitz as well.
Proof.
By Corollary 4.4, there are homeomorphisms $f_{i}:B(0,\varrho)\twoheadrightarrow B(0,\varrho)$, $i=1,\ldots,k$, and balls $\mkern 1.7mu\overline{\mkern-1.7muU\mkern-1.5mu}\mkern 1.5mu_{i}\subset B(0,\varrho)$ such that
(4.3) $F=f_{1}\circ\ldots\circ f_{k},\quad f_{i}|_{U_{i}}=\operatorname{id}.$
Let $\lambda\in(\delta,\varrho)$ be such that $\bigcup_{i=1}^{k}\mkern 1.7mu\overline{\mkern-1.7muU\mkern-1.5mu}\mkern 1.5mu_{i}\subset B(0,\lambda)$. Since $f_{i}$ are surjective homeomorphisms of $B(0,\varrho)$, we have $|f_{i}(x)|\to\varrho$ as $|x|\to\varrho$, so there is $\tau\in(\lambda,\varrho)$ such that
(4.4) $|f_{k}(x)|>\lambda,\quad|(f_{k-1}\circ f_{k})(x)|>\lambda,\ldots,|(f_{1}\circ\ldots\circ f_{k})(x)|>\lambda\quad\text{whenever }|x|>\tau.$
Let $\psi_{i}:B(0,\varrho)\twoheadrightarrow B(0,\varrho)$ be a diffeomorphism that maps $B(0,\delta)$ onto $U_{i}$ and equals identity on $B(0,\varrho)\setminus B(0,\lambda)$. Then $\psi_{i}^{-1}\circ f_{i}\circ\psi_{i}:B(0,\varrho)\twoheadrightarrow B(0,\varrho)$ is a homeomorphism which coincides with identity on $B(0,\delta)$. We set
$\widehat{F}:=\psi_{1}^{-1}\circ f_{1}\circ\psi_{1}\,\circ\,\psi^{-1}_{2}\circ f_{2}\circ\psi_{2}\,\circ\,\ldots\circ\,\psi_{k}^{-1}\circ f_{k}\circ\psi_{k}.$
Clearly, $\widehat{F}$ satisfies the second condition in (4.2), and (4.4) along with a simple induction imply that $\widehat{F}$ satisfies the first condition in (4.2). Moreover, if $F$ is locally bi-Lipschitz, we can assume that $f_{i}$ are locally bi-Lipschitz and hence $\widehat{F}$ is locally bi-Lipschitz. ∎
In accordance with our definitions of flat topological and bi-Lipschitz balls, we say that $S\subset\mathbb{R}^{n}$ is a flat topological (or flat bi-Lipschitz) sphere, if there is a (bi-Lipschitz) homeomorphism $F:\mathbb{S}^{n-1}\times(-1,1)\to\mathbb{R}^{n}$ such that $F(\mathbb{S}^{n-1}\times\{0\})=S$.
Remark 4.6.
There is an elegant characterization of flat topological (or flat bi-Lipschitz) spheres and balls in $\mathbb{R}^{n}$.
  • a)
    A topological (or bi-Lipschitz) sphere $S\subset\mathbb{R}^{n}$ is flat if and only if it is a (bi-Lipschitz) submanifold of $\mathbb{R}^{n}$, that is every point $x\in S$ has a neighborhood $U$ such that the pair $(U,S\cap U)$ is (bi-Lipschitz) homeomorphic to ${(\mathbb{B}^{n},\{(x_{1},\ldots,x_{n})\in\mathbb{B}^{n}}\,\,:\,\,x_{n}=0\})$.
  • b)
    A topological (or bi-Lipschitz) closed ball $D\subset\mathbb{R}^{n}$ is flat if and only if $\partial D$ is a flat topological (or flat bi-Lipschitz) sphere.
Obviously, a flat sphere is a submanifold, both in the topological and the bi-Lipschitz case. The reverse assertion in a) follows from the seminal paper of Brown [2, Lemma 6 and Theorem 4] (note that a flat sphere as per our definition is a bi-collared sphere in Brown’s paper), see also [5]. The bi-Lipschitz case is covered in [16, Theorem 7.8]. Assertion b) follows from the generalized Schönflies theorem proved by Brown [1, Theorem 5]; the bi-Lipschitz case again follows immediately from [16, Theorem 7.8].
Lemma 4.7 (Annulus theorem).
Let $S_{1},S_{2}\subset\mathbb{R}^{n}$ be two disjoint flat topological (or flat bi-Lipschitz) spheres such that $S_{1}$ is contained in the bounded component of $\mathbb{R}^{n}\setminus S_{2}$. Then the set bounded by $S_{1}$ and $S_{2}$ is (bi-Lipschitz) homeomorphic to the annulus $\mkern 1.7mu\overline{\mkern-1.7mu\mathbb{B}\mkern-1.5mu}\mkern 1.5mu^{n}\setminus B^{n}(0,1/2)$.
Brown and Gluck [3], proved that the annulus theorem (in the case of flat topological spheres) is a direct consequence of the stable homeomorphism theorem (in the case of homeomorphisms) and the generalized Schönflies theorem, [1, Theorem 5]. See Corollary on page 8 and Theorem 3.5(i) in [3]. Their argument is elementary and geometric (but very tricky) and it verbatim applies to the bi-Lipschitz case: the bi-Lipschitz stable homeomorphism theorem implies the bi-Lipschitz annulus theorem, see [27, Theorem 3.12].
Remark 4.8.
By a radial extension of a homeomorphism $f:\partial B(0,\varrho)\twoheadrightarrow\partial B(0,\varrho)$ we mean a mapping $\tilde{f}:\mathbb{R}^{n}\twoheadrightarrow\mathbb{R}^{n}$ defined as $\tilde{f}(x)=\varrho^{-1}|x|f(\varrho x/|x|)$ for $x\neq 0$ and $\tilde{f}(0)=0$. Clearly, $\tilde{f}$ is a homeomorphism and if $f$ is bi-Lipschitz, then $\tilde{f}$ is bi-Lipschitz.
The next result, [27, Theorem 3.16], is a generalization of Lemma 2.1 to the case of homeomorphisms and bi-Lipschitz maps. It will allow us to use Palais’ argument (Lemma 2.2) beyond the realm of diffeomorphisms (Lemma 5.1).
Lemma 4.9.
Let $F:\mkern 1.7mu\overline{\mkern-1.7muB\mkern-1.5mu}\mkern 1.5mu(0,\varrho)\to\mathbb{R}^{n}$, $F(0)=0$, be an orientation preserving (bi-Lipschitz) homeomorphism. Then, there is $\delta\in(0,\varrho/2)$ and a (bi-Lipschitz) homeomorphism $\widetilde{F}:\mkern 1.7mu\overline{\mkern-1.7muB\mkern-1.5mu}\mkern 1.5mu(0,\varrho)\to\mathbb{R}^{n}$ such that $\widetilde{F}=F$ on $\mkern 1.7mu\overline{\mkern-1.7muB\mkern-1.5mu}\mkern 1.5mu(0,\varrho)\setminus B(0,\varrho/2)$ and $F=\operatorname{id}$ on $\mkern 1.7mu\overline{\mkern-1.7muB\mkern-1.5mu}\mkern 1.5mu(0,\delta)$.
Proof.
Firstly, note that since $F(0)=0$, there is a $\delta\in(0,\varrho/4)$ for which $\mkern 1.7mu\overline{\mkern-1.7muB\mkern-1.5mu}\mkern 1.5mu(0,2\delta)\subset F(B(0,\varrho/2))$.
The set $\partial F(B(0,\varrho/2))=F(\partial B(0,\varrho/2))$ is a flat topological sphere; let $A$ denote the compact region between $\partial F(B(0,\varrho/2))$ and $\partial B(0,2\delta)$. By Lemma 4.7, there is a homeomorphism $H:\mkern 1.7mu\overline{\mkern-1.7muB\mkern-1.5mu}\mkern 1.5mu(0,\varrho/2)\setminus B(0,2\delta)\twoheadrightarrow A$ and therefore
$H^{-1}\circ F:\partial B(0,\varrho/2)\twoheadrightarrow\partial B(0,\varrho/2)\text{ and }H:\partial B(0,2\delta)\twoheadrightarrow\partial B(0,2\delta).$
We can find radial extensions of $H^{-1}\circ F$ and $H$ to homeomorphisms $\Phi:\mkern 1.7mu\overline{\mkern-1.7muB\mkern-1.5mu}\mkern 1.5mu(0,\varrho/2)\twoheadrightarrow\mkern 1.7mu\overline{\mkern-1.7muB\mkern-1.5mu}\mkern 1.5mu(0,\varrho/2)$ and $\Psi:\mkern 1.7mu\overline{\mkern-1.7muB\mkern-1.5mu}\mkern 1.5mu(0,2\delta)\twoheadrightarrow\mkern 1.7mu\overline{\mkern-1.7muB\mkern-1.5mu}\mkern 1.5mu(0,2\delta)$, respectively. By Lemma 4.5, we find homeomorphisms $\widetilde{\Phi}:\mkern 1.7mu\overline{\mkern-1.7muB\mkern-1.5mu}\mkern 1.5mu(0,\varrho/2)\twoheadrightarrow\mkern 1.7mu\overline{\mkern-1.7muB\mkern-1.5mu}\mkern 1.5mu(0,\varrho/2)$ and $\widetilde{\Psi}:B(0,2\delta)\twoheadrightarrow B(0,2\delta)$ such that $\widetilde{\Phi}=\Phi$ near $\partial B(0,\varrho/2)$ and $\widetilde{\Phi}=\operatorname{id}$ on $B(0,2\delta)$ and $\widetilde{\Psi}=\Psi$ near $\partial B(0,2\delta)$ and $\widetilde{\Psi}=\operatorname{id}$ on $B(0,\delta)$. In particular, $\widetilde{\Phi}=H^{-1}\circ F$ on $\partial B(0,\varrho/2)$ and $\widetilde{\Psi}=H$ on $\partial B(0,2\delta)$.
Eventually, we set
$\widetilde{F}(x)=\begin{cases}F(x)&\text{ for }x\in\mkern 1.7mu\overline{\mkern-1.7muB\mkern-1.5mu}\mkern 1.5mu(0,\varrho)\setminus\mkern 1.7mu\overline{\mkern-1.7muB\mkern-1.5mu}\mkern 1.5mu(0,\varrho/2),\\ H\circ\widetilde{\Phi}(x)&\text{ for }x\in\mkern 1.7mu\overline{\mkern-1.7muB\mkern-1.5mu}\mkern 1.5mu(0,\varrho/2)\setminus\mkern 1.7mu\overline{\mkern-1.7muB\mkern-1.5mu}\mkern 1.5mu(0,2\delta),\\ \widetilde{\Psi}(x)&\text{ for }x\in\mkern 1.7mu\overline{\mkern-1.7muB\mkern-1.5mu}\mkern 1.5mu(0,2\delta).\end{cases}$
Since $\widetilde{F}(\mkern 1.7mu\overline{\mkern-1.7muB\mkern-1.5mu}\mkern 1.5mu(0,\varrho/2)\setminus B(0,2\delta))=A$ and $H\circ\widetilde{\Phi}$ agrees with $F$ on $\partial B(0,\varrho/2)$, and $\widetilde{\Psi}=H=H\circ\widetilde{\Phi}$ on $\partial B(0,2\delta)$, $\widetilde{F}$ is indeed a homeomorphism. Clearly, it satisfies the required properties.
If $F$ is additionally assumed to be bi-Lipschitz, then by Lemma 4.7, $H$ is a bi-Lipschitz homeomorphism and so are the radial extensions $\Phi$ and $\Psi$. Therefore, by Lemma 4.5, $\widetilde{\Phi}$ and $\widetilde{\Psi}$ are bi-Lipschitz and so is $\widetilde{F}$. ∎
5. Proof of Theorems 1.6 and 1.8
Lemma 5.1.
Suppose that $H:\mkern 1.7mu\overline{\mkern-1.7muB\mkern-1.5mu}\mkern 1.5mu(0,\varrho)\to\mathbb{R}^{n}$, $H(0)=0$, is an orientation preserving (bi-Lipschitz) homeomorphism which can be extended as a (bi-Lipschitz) homeomorphism on a neighborhood of $\mkern 1.7mu\overline{\mkern-1.7muB\mkern-1.5mu}\mkern 1.5mu(0,\varrho)$. Then, for any $\varepsilon>0$, there is a (bi-Lipschitz) homeomorphism $\widetilde{H}:\mathbb{R}^{n}\twoheadrightarrow\mathbb{R}^{n}$ such that
(5.1) $\widetilde{H}(x)=\begin{cases}H(x)&\text{ if }x\in\mkern 1.7mu\overline{\mkern-1.7muB\mkern-1.5mu}\mkern 1.5mu(0,\varrho),\\ x&\text{ if }\operatorname{dist}(x,A)\geq\varepsilon,\end{cases}$
where $A=\mkern 1.7mu\overline{\mkern-1.7muB\mkern-1.5mu}\mkern 1.5mu(0,\varrho)\cup H(\mkern 1.7mu\overline{\mkern-1.7muB\mkern-1.5mu}\mkern 1.5mu(0,\varrho))$.
Proof.
The proof of this lemma is a verbatim copy of Lemma 2.2 with ‘homeomorphism’ or ‘bi-Lipschitz homeomorphism’ instead of ‘diffeomorphism’. Also, instead of Lemma 2.1, we use Lemma 4.9. ∎
Remark 5.2.
If $G:\mkern 1.7mu\overline{\mkern-1.7muB\mkern-1.5mu}\mkern 1.5mu(a,r)\twoheadrightarrow\mkern 1.7mu\overline{\mkern-1.7muB\mkern-1.5mu}\mkern 1.5mu(a,r)$ is a (bi-Lipschitz) homeomorphism, then it can be extended as a (bi-Lipschitz) homeomorphism on a neighborhood of $\mkern 1.7mu\overline{\mkern-1.7muB\mkern-1.5mu}\mkern 1.5mu(a,r)$. Indeed, it suffices to extend $G$ radially to the exterior of $\mkern 1.7mu\overline{\mkern-1.7muB\mkern-1.5mu}\mkern 1.5mu(a,r)$, cf. Remark 4.8.
Corollary 5.3.
Let $G:\mkern 1.7mu\overline{\mkern-1.7muB\mkern-1.5mu}\mkern 1.5mu(a,r)\twoheadrightarrow\mkern 1.7mu\overline{\mkern-1.7muB\mkern-1.5mu}\mkern 1.5mu(a,r)$ be an orientation preserving (bi-Lipschitz) homeomorphism. Then for any $\varepsilon>0$, there is a (bi-Lipschitz) homeomorphism $\widetilde{G}:\mathbb{R}^{n}\twoheadrightarrow\mathbb{R}^{n}$ such that $\widetilde{G}=G$ on $\mkern 1.7mu\overline{\mkern-1.7muB\mkern-1.5mu}\mkern 1.5mu(a,r)$ and $\widetilde{G}=\operatorname{id}$ outside $B(a,r+\varepsilon)$.
Proof.
The proof is the same as the proof of Corollary 2.4 when ‘diffeomorphism’ is replaced with ‘(bi-Lipschitz) homeomorphism’. Also, we use Lemma 5.1 in place of Lemma 2.2. To do so, one needs to check that $G_{1}$ (defined in the proof of Corollary 2.4) can be extended as a (bi-Lipschitz) homeomorphism on a neighborhood of $\mkern 1.7mu\overline{\mkern-1.7muB\mkern-1.5mu}\mkern 1.5mu(0,r)$. This is true due to Remark 5.2. ∎
Next, we proceed with the proof of Theorems 1.6 and 1.8 following closely that of Theorem 1.1.
Lemma 5.4.
Let $\{D_{i}\}_{i=1}^{\ell}$, $D_{i}\subset\mathbb{R}^{n}$, be a family of pairwise disjoint flat topological (or flat bi-Lipschitz) closed balls and let $p_{i}\in\mathring{D}_{i}$, $i=1,\ldots,\ell$, be given. Then there is an $\varepsilon_{o}>0$ such that for any $\varepsilon\in(0,\varepsilon_{o})$ there is a (bi-Lipschitz) homeomorphism $F_{\varepsilon}:\mathbb{R}^{n}\twoheadrightarrow\mathbb{R}^{n}$ satisfying $F_{\varepsilon}(\mkern 1.7mu\overline{\mkern-1.7muB\mkern-1.5mu}\mkern 1.5mu(p_{i},\varepsilon))=D_{i}$ for $i=1,\ldots,\ell$ and
$F_{\varepsilon}(x)=x\text{ if }\operatorname{dist}\Big{(}x,\bigcup\nolimits_{i=1}^{\ell}D_{i}\Big{)}\geq\varepsilon.$
Proof.
The proof is the same as the proof of Lemma 3.1 with ‘(bi-Lipschitz) homeomorphism’ in place of ‘diffeomorphism’. Instead of Lemma 2.2, we use Lemma 5.1. ∎
As in Section 3, combining Lemma 3.2 with Lemma 5.4 yields the Euclidean version of Theorems 1.6 and 1.8.
Corollary 5.5.
Let $U\subset\mathbb{R}^{n}$ be a domain. Suppose that $\{D_{i}\}_{i=1}^{\ell}$ and $\{D^{\prime}_{i}\}_{i=1}^{\ell}$, $D_{i},D^{\prime}_{i}\subset U$, are two families of pairwise disjoint flat topological (or flat bi-Lipschitz) closed balls. If $F_{i}:D_{i}\twoheadrightarrow D^{\prime}_{i}$, $i=1,2,\ldots,\ell$, are orientation preserving (bi-Lipschitz) homeomorphisms, then there is a (bi-Lipschitz) homeomorphism $F:\mathbb{R}^{n}\twoheadrightarrow\mathbb{R}^{n}$ such that $F|_{D_{i}}=F_{i}$ and $F=\operatorname{id}$ in $\mathbb{R}^{n}\setminus U$.
Proof.
The proof follows the proof of Corollary 3.3. One needs to write ‘(bi-Lipschitz) homeomorphism’ instead of ‘diffeomorphism’ and invoke Lemma 5.4 in place of Lemma 3.1 and Corollary 5.3 in place of Corollary 2.4. ∎
Proof of Corollary 1.9.
The proof is the same as the proof of Corollary 1.4 except for the fact that it uses Corollary 5.5 in place of Corollary 3.3. ∎
Proof of Theorems 1.6 and 1.8.
Like in the case of Theorem 1.1, we prove these results by reducing them to the Euclidean setting, i. e., to the case treated in Corollary 5.5.
Let $U$ be a domain containing $D_{i},D^{\prime}_{i}$ for $i=1,\ldots,\ell$ (note that we include the case $U=\mathcal{M}^{n}$). Let $K\subset U$ be a flat topological (or flat bi-Lipschitz) closed ball. Then, there are (bi-Lipschitz) homeomorphisms $H,H^{\prime}:\mathcal{M}^{n}\twoheadrightarrow\mathcal{M}^{n}$ such that $H(D_{i})\subset\mathring{K}$, $H^{\prime}(D_{i}^{\prime})\subset\mathring{K}$ for all $i=1,2,\ldots,\ell$, and $H$ and $H^{\prime}$ are identity outside $U$.
Below, we sketch the construction of $H$, $H^{\prime}$ is constructed in the same way. We have to modify the construction of $H$ from Theorem 1.1 since now we clearly cannot directly use a $1$-parameter group of diffeomorphisms.
For each $i=1,2,\ldots,\ell$, we create a chain of coordinate systems $\{(\Phi_{ij},U_{ij})\}_{j=1}^{N_{i}}$ in $U$ that connects $D_{i}$ to $\mathring{K}$, i. e.,
$\Phi_{ij}:U_{ij}\twoheadrightarrow\mathbb{B}^{n},\quad\mkern 1.7mu\overline{\mkern-1.7muU\mkern-1.5mu}\mkern 1.5mu_{ij}\subset U,\quad D_{i}\subset U_{i1},\quad U_{ij}\cap U_{i,j+1}\neq\varnothing,\quad U_{iN_{i}}\cap\mathring{K}\neq\varnothing,$
and $\Phi_{ij}$ is a (bi-Lipschitz) homeomorphism.
We can also guarantee that the chains corresponding to different balls are disjoint i.e.,
$\Big{(}\bigcup\nolimits_{j=1}^{N_{s}}U_{sj}\Big{)}\cap\Big{(}\bigcup\nolimits_{j=1}^{N_{t}}U_{tj}\Big{)}=\varnothing\qquad\text{whenever }s\neq t.$
Now for each $i$, we construct a (bi-Lipschitz) homeomorphism $H_{i}:\mathcal{M}^{n}\twoheadrightarrow\mathcal{M}^{n}$ that is identity outside $\bigcup_{j=1}^{N_{i}}U_{ij}$ and such that $H_{i}(D_{i})\subset\mathring{K}$. We construct $H_{i}$ as a composition of (bi-Lipschitz) homeomorphisms $H_{i}=H_{iN_{i}}\circ\ldots\circ H_{i1}$ such that
$H_{i1}(D_{i})\subset U_{i2},\quad H_{i2}(H_{i1}(D_{i}))\subset U_{i3},\ \ldots,\ H_{iN_{i}}(\ldots(H_{i2}(H_{i1}(D_{i}))))\subset\mathring{K}.$
Each such a homeomorphism $H_{ij}$ can be constructed in the coordinate system $(\Phi_{ij},U_{ij})$ using a $1$-parameter group of diffeomorphisms. Note that the homeomorphisms $H_{i}$ glue to a (bi-Lipschitz) homeomorphism $H:\mathcal{M}^{n}\twoheadrightarrow\mathcal{M}^{n}$ with required properties.
The remaining part of the proof is the same as in the proof of Theorem 1.1 except that we use Corollary 5.5 in place of Corollary 3.3. ∎